OpenStudy (anonymous):

convert 30 cm/s to m/h

OpenStudy (anonymous):

1 cm = 10^(-2)m 1 s = (60)^(-2)h

OpenStudy (anonymous):

Do u know that?

hero (hero):

Are we converting centimeters per second to miles per hour or meters per hour?

OpenStudy (kropot72):

2.54 cm = 1 inch 30 cm/s = \[\frac{30}{2.54}inches/s\] At that speed in inches per second the distance travelled in one hour =\[\frac{30}{2.54}\times 60\times 60inches\] There are 5280 * 12 inches in a mile. Therefore speed in miles per hour is\[\frac{30\times 60\times 60}{2.54\times 5280\times 12}=?mph\]

OpenStudy (unklerhaukus):

miles is \([\text {mi}]\) meters is \([\text {m}]\)

OpenStudy (unklerhaukus):

\[30 \frac{[\text {cm}]}{[\text s]}\]\[=30 \frac{[\text {cm}]}{[\text s]}\times\frac{[\text m]}{100[\text {cm}]}\times\frac{3600[\text s]}{[\text{hr}]}\]\[=\frac{30\times3600}{100} \frac{\cancel{[\text {cm}]}}{\cancel{[\text s]}}\times\frac{[\text m]}{\cancel{[\text {cm}}]}\times\frac{\cancel{[\text s]}}{[\text{hr}]}\]\[=30\times36\frac{[\text m]}{[\text{hr}]}\]

hero (hero):

If miles is mi, then what is mph?

hero (hero):

meters per hour I suppose.

OpenStudy (kropot72):

@UnkleRhaukus and @Hero Point taken :0

hero (hero):

I was asking because if it was meters, then it is an easy solve: 30 cm/s = x m/h 30/100s = x/h (30h)/(100s) = x (30*3600s)/100s = x (30*3600)/100 = x (30*36) = x 1080 = x